OpenStudy (anonymous):

The department store has recently raised the price of a television set to $154.76. this was a raise of 6%. what was the price of a television set before the raise?

OpenStudy (anonymous):

Let the price before the raise be : x x + 6% of x = 154.76 \[x + 6x/100 = 154.76\] \[106x/100 = 154.76\] 106x = 15476 Can you now find the value of x??

OpenStudy (anonymous):

thank you i found the answer:)

OpenStudy (anonymous):

Welcome dear..